OpenStudy (anonymous):

What value of n solves the equation? 4^n = 1/64 A. –3 B. 3 C. 16 D. 256

OpenStudy (michele_laino):

please note that: \[4^{n}=(2^{2})^{n}=2^{2n}\] and \[\frac{ 1 }{ 64 }=64^{-1}=(2^{6})^{-1}=2^{-6}\] so please substitute the above equalities into your equation

OpenStudy (anonymous):

I still dont get it, we just started this in class today but im still really confused by all of it.

OpenStudy (michele_laino):

please do you understand the equalities that I wrote above?

OpenStudy (anonymous):

Not really i asked my sister if she could help me but she said she doesnt understand it either :/ im not the best at math.

OpenStudy (michele_laino):

ok! Im going to explaint them you now!

OpenStudy (anonymous):

Thank you!

OpenStudy (michele_laino):

we have \[\frac{ 1 }{ 64 }\] now in that fraction the exponent of 64 is 1, do you agree?

OpenStudy (anonymous):

But wouldnt i be -1? My teacher said when you have a negative exponent it turns into a fraction.

OpenStudy (anonymous):

it*

OpenStudy (michele_laino):

I know what you say, nevertheless we have : \[(64)^{1}=64\] do you agree?

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

ok! now there is a property of powers which is: \[\frac{ 1 }{ (a)^{m} }=a ^{-m}\] please be careful, that above is a property do you agree?

OpenStudy (anonymous):

This is the part that im confused on, but i think i understand. I cant remember what the (a) is. But i think its a property

OpenStudy (michele_laino):

in our casgeneral e acan be any real positive number, whereas n can be any unteger positive number. Now in our case a=64 and m=1

OpenStudy (michele_laino):

sorry: in general case a can be any real positive number....

OpenStudy (anonymous):

ohh okay that would mean 1/a^1 correct?

OpenStudy (michele_laino):

\[\frac{ 1 }{ a ^{1} }\] yes it is correct

OpenStudy (anonymous):

Okay i understand that now

OpenStudy (michele_laino):

so if we apply our property, we can write: \[\frac{ 1 }{ 64 }=(64)^{-1}\] do you agree?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

now, please note that: \[64=4*4*4=2^{2}*2^{2}*2^{2}=2^{6}\] because, ise multiplication of powers with equal base is a power whose base is the same base and the exponent is the sum of the exponents of the powers which we want to multiply

OpenStudy (isaiah.feynman):

-3. Due to multiplication tables.

OpenStudy (anonymous):

Okay i understand that part

OpenStudy (isaiah.feynman):